# Effective synchronization amid noise-induced chaos

### v2

@@ VI Conclusion
The mechanism we reported here is hardly a compelling explanation of neuronal information transfer and synchronized action among autonomous agents. Nonetheless, this mechanism shows an implicit organization, i.e., synchronization, hidden in apparent chaos. Thus, it adds to the possible ways that evolved systems such as biological organisms might exploit complex dynamics to generate subtle, adaptive behavior. It seems worthwhile to explore such mechanisms further.
@@ Acknowledgements.
We benefited from discussions with Efi Efrati and Colin Scheibner. We thank Chuanyi Wang for providing the efficient functional form of the quintic maps. We also thank Gil Ariel, Eli Barkai, Omer Granek, Andrej Košmrlj, Adilson Motter, Muhittin Mungan, Colin Scheibner, and Tony Song for helpful comments on the manuscript. B.S. acknowledges support from the Princeton Center for Theoretical Science. B.S. would like to thank the members of the James Frank Institute of the University of Chicago for their warm hospitality, where the research was initiated.
Appendix A Materials and Methods
A.1 Binless estimation methods using samples
@@
$q_{{\mathrm{A}}n}^{(k)}\equiv q^{(k)}_{\mathrm{A}}\left(\varphi=\frac{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}+\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}}{2}\right)=\frac{1}{N}\frac{1}{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}-\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}}$ (8)
@@
(and the same for $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$), where $\gamma=0.577$ is the Euler-Mascheroni constant. See Fig. 3 and SI Appendix, Fig. S6 for a typical evolution of the entropy and KLD for $\Lambda=0.141$ and $\Lambda=-0.141$, respectively.
@@
The loss of samples to machine precision is a prevalent limitation. In the long realization ($k=10^{4}$) of Sec. IV, $407$ and $439$ samples remained in the end for ensembles $\mathrm{A}$} and $\mathrm{B}$}, respectively, which means that our estimates are still statistically meaningful. On the other hand, this implies that the low-entropy distributions (which are those that usually culminate in loss of samples) did not appear much; see Fig. 5, where we see no distribution with S<21S<-21. Upon continuing this simulation, there were no longer any distinguishable phase samples past $k=1.2\cdot 10^{4}$, at which point the dynamics have attained entropies between 34<S<21-34<S<-21. Were lower-entropy distributions to appear, our results of the main text would improve further, as they are particularly useful for having a small discrepancy $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$.
A.2 Identifying a fiducial phase
@@
Supplementary Material: Effective synchronization amid noise-induced chaos
Appendix S1 Instantaneous fiducial phase
Figure S1 illustrates a phase probability distribution defined in Sec. II of the main text. It pictures a time interval long after the most recent kick, but before the next kick. It shows how the phase distribution $q(\varphi)$ and the fiducial phase $\varphi_{\mathrm{f}}$ move in time.
Appendix S2 Example of a phase map
In Sec. II of the main text, we defined a “phase map” $\psi(\varphi)$ which encapsulates the effect of a kick on a particular nonlinear oscillator. Here, we illustrate how this phase map is determined for a specific dynamical system and type of kick.
In this example, we construct the phase map for perturbations around the limit cycle of a double-well Stuart-Landau dynamics [28, 29] in the complex plane $z$. Here, $z\equiv x+iy$ evolves according to
$\frac{dz}{dt}=\frac{1}{2}\left[f(z+0.4+0.4i,5,2.75)+f(z-0.4-0.4i,-1,-4)\right],$ (S1a)
where
$f(z,a,b)=(1+ai)z-(1+bi)|z|^{2}z.$ (S1b)
The numbers were chosen such that a stable periodic orbit would be observed; see full black oval in Fig. S2. The period is $T\simeq 1.98$. We designate some point along the limit-cycle orbit as the phase origin, shown as an open square in Fig. S2. We define the phase position $\varphi$ of given point on the orbit as the time required to move from the phase origin to the given point, relative to the period $T$ (the black ticks along the black oval in Fig. S2). Thus, at the phase position $\varphi=1$, the given point has traversed the whole cycle and returned to the phase origin.
We now consider the effect of two particular kicks, defined to be a displacement of $z$ by an amount $0.15+0i$ in Fig. S2(a) and $0.45+0i$ in Fig. S2(b). We have chosen this displacement so that any point $\varphi$ on the limit cycle returns to it virtually completely in a time $9T$ or less. Thus after any number of cycles greater than $9$, the final phase $\psi$ of the oscillator is constant. Displacements bigger than $\simeq 0.46$ cause some of the initial phases to escape the periodic orbit’s basin of attraction. In that sense, the latter displacement (b) is a very strong forcing; the former (a) is much milder.
If one kicks an oscillator that was at $\varphi=0$ (empty squares), it will then undergo the trajectories shown with the dashed red curves in Fig. S2. After $9T,10T,11T,\ldots$, we find, to good accuracy, that it reaches the points marked by solid squares, which is found to have a phase $\psi\simeq-0.13$ for the weak forcing (a) and $\psi\simeq-0.44$ for the strong forcing (b). (As before, $\psi$ is still defined $\mathrm{mod}\,1$. However, so to obtain the smooth collection of blue empty points in Fig. S3, we determined integer offsets by counting the number of cycles completed during $9T$ relative to the $9$ cycles that an unperturbed oscillator would have completed during $9T$.) Likewise, kicking oscillators that are at $\varphi\simeq 0.35$ (empty triangles) and $\varphi\simeq 0.6$ (empty diamonds), through the dotted-orange and full-blue trajectories, they will, respectively, reach $\psi\simeq 0.73$ and $\psi\simeq 0.36$ for the weak forcing (a) and $\psi\simeq 1.37$ and $\psi\simeq-0.66$ for the strong forcing (b), which are depicted with solid triangles and diamonds. We remind that were the three not kicked, after $9T$ they would have returned back to their original positions (0, $0.35$, and $0.6$), and that any phases that differ by an integer necessarily correspond to the same point on the orbit.
To determine the final phase $\psi$ for arbitrary initial phases $\varphi$, we repeat this process for many closely spaced initial $\varphi$’s. The result is the phase maps shown in Fig. S3, which depend on the displacement (forcing magnitude), shown in Fig. S3. The lags between the initial and final phases are shown with the empty blue symbols in Fig. S2. The squares, triangles, and diamonds correspond to the final phase $\psi$ of $\varphi=0$, $\varphi\simeq 0.35$, and $\varphi\simeq 0.6$, respectively, under each forcing. Constructing the phase map $\psi(\varphi)$ amounts to taking $\mathrm{mod}\,1$ of the resulting lags. Since the system is rotation invariant (as the next kick is also randomly timed according to a uniform distribution $[0,1)$), we shift all phases such that, arbitrarily, $\psi(\varphi=0)=0$. This allows the comparison with the cubic phase map considered in the main text. Upon this shift and taking the moduli, we find the phase maps depicted with connected red points. Indeed the very strong forcing $z\to z+0.45$ leads to a very violent phase map (even having an apparent discontinuity in the spreading factor $d\psi/d\varphi$ at $\varphi\simeq 0.32$). The mild forcing $z\to z+0.15$, which is the limit with which we are concerned in this work, has given rise to a well-behaved and smooth phase map with a single minimum and maximum. As we explained in the main text, broad minima and maxima are the ones facilitating the formation of sharply-peaked and thence low-entropy distributions. This suggests that the findings of the main text for the simplistic cubic phase map we chose should apply in ‘real-world’ forcings as well.
Appendix S3 Phase map iteration
In Fig. S4 we illustrate the changes in a set of initial phases $\{\varphi_{n}\}$ through an iteration of the phase map of Eq. (2) of the main text.
Appendix S4 Dependence on stochastic realization
In Fig. 2 of the main text, we show a sequence of distributions obtained for two agents under the influence of common noise. Despite the different initial conditions, for $\Lambda<0$ we observe synchronization, whereas for $\Lambda>0$ we observe convergence of ensembles. To emphasize that a common noise is a necessary condition, in Fig. S5 we show two ensembles, starting from the same distribution, which are subjected to different noises. Evidently, since the information about the ambient noise is not shared, the distributions do not synchronize with $\Lambda<0$ and do not converge for $\Lambda>0$.
Appendix S5 Synchronization for $\Lambda<0$
Here we revisit the discussion of Sec. III of the main text for a synchronizing phase map with $\Lambda<0$. That is, the agents’ phases agree up to a gradually-diminishing uncertainty. For $\Lambda<0$, both distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ and $q^{(k)}_{\mathrm{B}}(\varphi)$ are expected to become a sharp peak centered around the same phase value. To test this, we initiate both distributions to be concentrated uniformly into eccentric narrow sectors of angular width $u=0.05$, one centered at $\varphi=0.025$ and the other at $\varphi=0.525$. While indeed both entropies decay to -$\infty$ (see Fig. S6(b)), both KLDs have reached a different nonzero plateau (see Fig. S6(a)).
[…]
$q^{(k)}_{\mathrm{A}}(\varphi)\simeq q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})\left|\left.\frac{d\psi}{d\varphi}\right|_{\hat{\varphi}+\beta^{(k)}}\right|^{-1},$ (S3)
$D^{(k)}(\mathrm{A}\|\mathrm{B})=\int_{0}^{1}d\hat{\varphi}q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})\ln\frac{q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})}{q_{\mathrm{B}}^{(k-1)}(\hat{\varphi})}=D^{(k-1)}(\mathrm{A}\|\mathrm{B}),$ (S4)
and similarly $D^{(k)}(\mathrm{B}\|\mathrm{A})=D^{(k-1)}(\mathrm{B}\|\mathrm{A})$. Thus, indeed the KLD is capable of remaining constant for two infinitely narrow distributions, as they either “stretch” or “contract” together to the same extent. Similar reasoning explains why, while the entropies decay to -$\infty$, the difference among them for these two narrow distributions remains constant for low entropy values in Fig. S6(b).
From similar consideriations, observe the $25\lesssim k\lesssim 65$ regime in Fig. 3(a) of the main text, where the dynamics of $\Lambda>0$ is shown. As seen in Fig. 2, $\Lambda=0.141$ panels of the main text, indeed the distributions are very narrow during these steps (which also manifests in a low entropy in Fig. 3(b) of the main text). This regime mimics the $\Lambda<0$ case at late times, where the narrow distributions change width together and barely mix, which manifests in a constant KLD in Fig. 3(a) of the main text. As we argued here, this nonzero KLD need not imply that these sharp distributions are not useful for communication, so long as they share means. However, since $\Lambda>0$, the distributions are bound to get sufficiently wide eventually, continue the mixing, and thus the KLD resumes its decay.
Appendix S6 Effect of the bandwidth

### v3

@@ VI Conclusion
In conclusion, our effective synchronization mechanism amid strong noise shows an implicit organization, i.e., synchronization, hidden in apparent chaos. It permits two or more agents to generate a stream of shared information that facilitates synchronization and thereby enables cooperative activity. It seems worthwhile to explore such mechanisms further.
@@ Acknowledgements.
We benefited from discussions with Charlie Edelson, Efi Efrati, Gautam Reddy, Colin Scheibner, and Chuanyi Wang. We also thank Gil Ariel, Eli Barkai, Omer Granek, Andrej Košmrlj, Adilson Motter, Muhittin Mungan, Colin Scheibner, and Tony Song for helpful comments on the manuscript. B.S. acknowledges support from the Princeton Center for Theoretical Science. B.S. would like to thank the members of the James Frank Institute of the University of Chicago for their warm hospitality, where this research was initiated.
Appendix A Example of a phase map
In Sec. II, we defined a “phase map” $\psi(\varphi)$ which encapsulates the effect of a kick on a particular nonlinear oscillator. Here, we illustrate how this phase map is determined for a specific dynamical system and type of kick.
In this example, we construct the phase map for perturbations around the limit cycle of a modifed Stuart-Landau dynamics in the complex plane $z$. Here, $z\equiv x+iy$ evolves according to
$\frac{dz}{dt}=\frac{1}{2}\bigl[$ $f(z+0.4+0.4i,5,2.75)$
$+f(z-0.4-0.4i,-1,-4)\bigr],$ (8a)
where
$f(z,a,b)=(1+ai)z-(1+bi)|z|^{2}z.$ (8b)
The conventional Stuart-Landau model [60, 61] exhibits either a stable fixed point or a harmonic limit cycle, characterized by a fixed radius and uniform angular velocity on the $z$ plane. Here, we aim to show that the phase-reduction procedure applies to oscillators with more complex limit cycles, particularly those which are not a circle in phase space nor have a uniform tangential velocity. The numbers in Eq. (8) were chosen such that a stable periodic orbit would be observed; see full black oval in Fig. 6. The period is $T\simeq 1.98$. We designate some point along the limit-cycle orbit as the phase origin, shown as an open square in Fig. 6. We define the phase position $\varphi$ of given point on the orbit as the time required to move from the phase origin to the given point, relative to the period $T$ (the black ticks along the black oval in Fig. 6). Thus, at the phase position $\varphi=1$, the given point has traversed the whole cycle and returned to the phase origin.
Figure 6: Trajectories from the dynamical system of Eq. (8). The limit cycle (period $T=1.98$) is shown in full black oval. The black ticks along it are equispaced $\varphi\in[0,1)$ values, separated by $\Delta t/T=0.05$. ($\varphi=0$ is located at $0+0.4i$, overlapping with the empty square.) We show the resulting trajectories from kicking oscillators positioned at $\varphi=0$ (empty square), $\varphi\simeq 0.35$ (empty triangle), and $\varphi\simeq 0.6$ (empty diamond) by (a) $z\to z+0.15$ and (b) $z\to z+0.45$. The first three small points on each trajectory are the positions of the oscillators after $t=0,(1/2)T,T$. After a time $9T$, the oscillators reach the stable orbit with new phase values, $\psi\simeq-0.13$ (full square), $\psi\simeq 0.73$ (full triangle), and $\varphi\simeq 0.36$ (full diamond) in (a), and $\psi\simeq-0.44$ (full square), $\psi\simeq 1.37$ (full triangle), and $\varphi\simeq-0.66$ (full diamond) in (b).
We now consider the effect of two particular kicks, defined to be a displacement of $z$ by an amount $0.15+0i$ in Fig. 6(a) and $0.45+0i$ in Fig. 6(b). We have chosen this displacement so that any point $\varphi$ on the limit cycle returns to it virtually completely in a time $9T$ or less. Thus after any number of cycles greater than $9$, the final phase $\psi$ of the oscillator is constant. Displacements bigger than $\simeq 0.46$ cause some of the initial phases to escape the periodic orbit’s basin of attraction into that of the stable fixed point at $z\simeq 1.5-0.56i$. In that sense, the latter displacement (b) is a very strong forcing; the former (a) is much milder.
If one kicks an oscillator that was at $\varphi=0$ (empty squares), it will then undergo the trajectories shown with the dashed red curves in Fig. 6. After $9T,10T,11T,\ldots$, we find, to good accuracy, that it reaches the pointx marked by solid squares, which is found to have a phase $\psi\simeq-0.13$ for the weak forcing (a) and $\psi\simeq-0.44$ for the strong forcing (b). (As before, $\psi$ is still defined $\mathrm{mod}\,1$. However, so to obtain the smooth collection of blue empty points in Fig. 7, we determined integer offsets by counting the number of cycles completed during $9T$ relative to the $9$ cycles that an unperturbed oscillator would have completed during $9T$.) Likewise, kicking oscillators that are at $\varphi\simeq 0.35$ (empty triangles) and $\varphi\simeq 0.6$ (empty diamonds), through the dotted-orange and full-blue trajectories, they will, respectively, reach $\psi\simeq 0.73$ and $\psi\simeq 0.36$ for the weak forcing (a) and $\psi\simeq 1.37$ and $\psi\simeq-0.66$ for the strong forcing (b), which are depicted with solid triangles and diamonds. We have picked $9T$ as our waiting time to guarantee that all oscillators have returns to the limit cycle within a small margin, so they can be once again uniquely described by a phase variable. We remind that were the three not kicked, after $9T$ they would have returned back to their original positions (0, $0.35$, and $0.6$), and that any phases that differ by an integer necessarily correspond to the same point on the orbit.
Figure 7: Phase lags from the dynamical system of Eq. (8). Empty blue symbols are phase shifts as seen in Fig. 6. The square, diamond, and triangle symbols correspond to the three trajectories in Fig. 6. Since only the fractional parts of phases are significant, each symbol was shifted upwards or downwards by an integer so as to obtain a continuous dependence in $\varphi$. In addition, this curve would be shifted vertically and horizontally by choosing a different point on the orbit as the phase origin. Red points show the corresponding phase map as we defined in the main text. Namely, the phase origin and integer-valued shifting were chosen so that $\psi(0)=0$. Further, each $\psi(\varphi)$ value was shifted to lie between 0 and 1.
To determine the final phase $\psi$ for arbitrary initial phases $\varphi$, we repeat this process for many closely spaced initial $\varphi$’s. The result is the phase maps shown in Fig. 7, which depend on the displacement (forcing magnitude), shown in Fig. 7. The lags between the initial and final phases are shown with the empty blue symbols in Fig. 6. The squares, triangles, and diamonds correspond to the final phase $\psi$ of $\varphi=0$, $\varphi\simeq 0.35$, and $\varphi\simeq 0.6$, respectively, under each forcing. Constructing the phase map $\psi(\varphi)$ amounts to taking $\mathrm{mod}\,1$ of the resulting lags. Since the system is rotation invariant (as the next kick is also randomly timed according to a uniform distribution $[0,1)$), we shift all phases such that, arbitrarily, $\psi(\varphi=0)=0$. This allows the comparison with the cubic phase map of Eq. (2). Upon this shift and taking the moduli, we find the phase maps depicted with connected red points. Indeed the very strong forcing $z\to z+0.45$ leads to a very violent phase map (even having an apparent discontinuity in the spreading factor $d\psi/d\varphi$ at $\varphi\simeq 0.32$). The mild forcing $z\to z+0.15$, which is the limit with which we are concerned in this work, has given rise to a well-behaved and smooth phase map with a single minimum and maximum. As we explained in Sec. II, broad minima and maxima are the ones facilitating the formation of sharply-peaked and thence low-entropy distributions. This suggests that our findings in Secs. III and IV for the simplistic cubic phase map we chose should apply in ‘real-world’ forcings as well.
Appendix B Binless estimation methods using samples
@@
$q_{{\mathrm{A}}n}^{(k)}$ $\equiv$ $q^{(k)}_{\mathrm{A}}\left(\varphi=\frac{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}+\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}}{2}\right)$ (9a)
$=$ $\frac{1}{N}\frac{1}{\tilde{\varphi}^{(k)}_{{\mathrm{A}}n+1}-\tilde{\varphi}^{(k)}_{{\mathrm{A}}n}}$ (9b)
@@
(and the same for $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$), where $\gamma=0.577$ is the Euler-Mascheroni constant. See Figs. 3 and 10 for a typical evolution of the entropy and KLD for $\Lambda=0.141$ and $\Lambda=-0.141$, respectively.
@@
The loss of samples to machine precision is a prevalent limitation. In the long realization ($k=10^{4}$) of Sec. IV, $407$ and $439$ samples remained in the end for ensembles $\mathrm{A}$} and $\mathrm{B}$}, respectively, which means that our estimates are still statistically meaningful. On the other hand, this implies that the low-entropy distributions (which are those that usually culminate in loss of samples) did not appear much; see Fig. 5, where we see no distribution with S<21S<-21. Upon continuing this simulation, there were no longer any distinguishable phase samples past $k=1.2\cdot 10^{4}$, at which point the dynamics have attained entropies between 34<S<21-34<S<-21. Were lower-entropy distributions to appear, our results in Sec. IV would improve further, as they are particularly useful for having a small discrepancy $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$.
Appendix C Phase map iteration
In Fig. 8 we illustrate the changes in a set of initial phases $\{\varphi_{n}\}$ through an iteration of the phase map of Eq. (2).
Figure 8: Two phase distributions $q^{(k)}(\varphi)$ before (blue full circles) and after (orange empty circles, with inflected axes) a kick, obtained for a typical realization of the cubic phase map (Eqs. (1) and (2)) with $A=9.32$ ($\Lambda=0.141$). The cubic phase map is drawn with a red line. One sees how the ‘sporadic’ region of the initial distribution in the vicinity the phase map’s extremum ($\varphi\simeq 0.7$) is condensed into a prominent peak at $\psi\simeq 0.3$. Conversely, the surroundings of the peak at $\varphi\simeq 0.95$ are widened at $\psi=0.7$, while its height decreased from $\mathcal{O}(10)$ to $\mathcal{O}(1)$. The phase distributions are sampled from $N=500$ oscillators.
Appendix D Dependence on stochastic realization
In Fig. 2, we show a sequence of distributions obtained for two agents under the influence of common noise. Despite the different initial conditions, for $\Lambda<0$ we observe synchronization, whereas for $\Lambda>0$ we observe convergence of ensembles. To emphasize that a common noise is a necessary condition, in Fig. 9 we show two ensembles, starting from the same distribution, which are subjected to different noises. Evidently, since the information about the ambient noise is not shared, the distributions do not synchronize with $\Lambda<0$ and do not converge for $\Lambda>0$.
$\Lambda=-0.141$ Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption $\Lambda=0.141$ Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
Figure 9: A sequence of two, initially-identical phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to different noises in a typical realization of Eqs. (1) and (2) with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\sim\mathrm{U}[0.5,0.55)$. The distributions are drawn on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. Since the noise is different for each ensemble, $\Lambda<0$ produces different synchronized states, whereas $\Lambda>0$ produces erratic distributions that do not converge unto each other. The distributions are sampled using $N=500$ oscillators via the nearest-neighbor distances; see Eq. (9b). The initial samples drawn for ensemble $\mathrm{A}$} are copied to ensemble $\mathrm{B}$}, so the starting point is truly identical.
Appendix E Below-threshold synchronization
Here we revisit the discussion of Sec. III for a synchronizing phase map with $\Lambda<0$. That is, the agents’ phases agree up to a gradually-diminishing uncertainty. For $\Lambda<0$, both distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ and $q^{(k)}_{\mathrm{B}}(\varphi)$ are expected to become a sharp peak centered around the same phase value. To test this, we initiate both distributions to be concentrated uniformly into eccentric narrow sectors of angular width $u=0.05$, one centered at $\varphi=0.025$ and the other at $\varphi=0.525$. While indeed both entropies decay to -$\infty$ (see Fig. 10(b)), both KLDs have reached a different nonzero plateau (see Fig. 10(a)).
Figure 10: (a) Kullback-Leibler divergences (KLDs) $D(\mathrm{A}\|\mathrm{B})$ and $D(\mathrm{B}\|\mathrm{A})$ among two distributions and (b) entropies $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ of each distribution, as obtained for the waiting time realization $\{\beta^{k}\}$ of Fig. 2 with $\Lambda=-0.141$. The number of phase samples is $N=500$. The initial distribution per ensemble is obtained with width $u=0.05$. The numerical errors in the estimation of the KLDs and entropies are $\Delta D=0.05$ and $\Delta S=0.2$, respectively.
[…]
$q^{(k)}_{\mathrm{A}}(\varphi)\simeq q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})\left|\left.\frac{d\psi}{d\varphi}\right|_{\hat{\varphi}+\beta^{(k)}}\right|^{-1},$ (12)
$D^{(k)}(\mathrm{A}\|\mathrm{B})$ $=$ $\int_{0}^{1}d\hat{\varphi}q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})\ln\frac{q_{\mathrm{A}}^{(k-1)}(\hat{\varphi})}{q_{\mathrm{B}}^{(k-1)}(\hat{\varphi})}$ (13)
$=$ $D^{(k-1)}(\mathrm{A}\|\mathrm{B}),$ (14)
and similarly $D^{(k)}(\mathrm{B}\|\mathrm{A})=D^{(k-1)}(\mathrm{B}\|\mathrm{A})$. Thus, indeed the KLD is capable of remaining constant for two infinitely narrow distributions, as they either “stretch” or “contract” together to the same extent. Similar reasoning explains why, while the entropies decay to -$\infty$, the difference among them for these two narrow distributions remains constant for low entropy values in Fig. 10(b).
From similar consideriations, observe the $25\lesssim k\lesssim 65$ regime in Fig. 3(a), where the dynamics of $\Lambda>0$ is shown. As seen in Fig. 2, $\Lambda=0.141$ panels, indeed the distributions are very narrow during these steps (which also manifests in a low entropy in Fig. 3(b)). This regime mimics the $\Lambda<0$ case at late times, where the narrow distributions change width together and barely mix, which manifests in a constant KLD in Fig. 3(a). As we argued here, this nonzero KLD need not imply that these sharp distributions are not useful for communication, so long as they share means. However, since $\Lambda>0$, the distributions are bound to get sufficiently wide eventually, continue the mixing, and thus the KLD resumes its decay.
Appendix F Instantaneous fiducial phase
Figure 11 illustrates a phase probability distribution defined in Sec. II. It pictures a time interval long after the most recent kick, but before the next kick. It shows how the phase distribution $q(\varphi)$ and the fiducial phase $\varphi_{\mathrm{f}}$ move in time.
Figure 11: Relation between the fiducial phase position $\varphi_{\mathrm{f}}$ and the instantaneous fiducial phase $\varphi_{\mathrm{f}}+0.2T$ at a later time $t=0.2T$, as measured by a given agent. Orange dotted curve represents the distribution $q(\varphi)$ for some particular noise history. The highest peak of this distribution is then $\varphi_{\mathrm{f}}$, and is marked with a dashed line. At time $t=0$ this is the actual phase distribution. At a time $t=0.2T$ later, every phase point, including the agent’s own oscilllator’s phase, has advanced through $0.2$ cycles. The probability distribution $q(\varphi-0.2T)$ and fiducial phase $\varphi_{\mathrm{f}}+0.2T$ at that moment is shown by the blue line.
Appendix G Identifying a fiducial phase
@@
Appendix H Effect of the bandwidth